# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

### v1

@@ 6.6. Ablation Studies
Table 5. Impact of three major stage in SensorChatC. Bold values highlight the best results.
@@
Impact of Sensor Features and Pretraining Loss Functions We next evaluate the impact of sensor features and loss functions during offline encoder pretraining. Specifically, we compare using statistical features (e.g., mean acceleration) versus raw time series inputs, and IMU2CLIP loss (Moon et al., 2023b) versus our proposed contrastive sensor-text pretraining loss for partial context (see Sec.4.2). For IMU2CLIP, text samples are generated by combining all valid labels into one sentence. We report the online querying accuracy and multiple-choice answer accuracy to assess the influence to sensor information extraction. As shown in Table 6, statistical features result in low querying and answer accuracies. This validates our motivation to design SensorChat that high-dimensional time series sensor data are critical for fine-grained activity information. IMU2CLIP training, even with fine-grained data, yields poorer querying and answering accuracies, highlighting its limited ability associating sensor embeddings from partial text query. Our proposed loss function, which aligns sensor and text encoders for partial context queries, proves more effective. These findings emphasize the importance of selecting appropriate sensor features and loss functions during pretraining in order to achieve high-performance QA.
@@
Table 7. Impact of various design choices for LLMs in SensorChatC. Bold values highlight the best results.
@@
Figure 10. Sensitivity of key hyperparameters.
6.7. Sensitivity Analyses
@@
Figure 11. Generalization of key learning components in SensorChat.
6.8. Generalization and Robustness
@@
7. Real User Study
We conduct a user study to evaluate SensorChat’s real-world applicability. In addition to dataset-based evaluations in Sec. 6 which focus on quantitative questions, this study places SensorChat into the “wilderness”. Participants will interact with SensorChat immediately after data collection and may pose arbitrary questions, especially the qualitative and open-ended ones. Our study is approved by the Institutional Review Board Committee.
User Study Setup We recruited eight volunteers (five males and three females) who were instructed to follow their normal daily routines while carrying a smartphone with the ExtraSensoryApp (Vaizman et al., 2018a). The smartphone models include Huawei Mate 10 Pro, LG G7 ThinQ and Google Pixel 2. The app automatically collected multimodal sensor data and transmitted it to SensorChat whenever a network connection was available. Reporting activity labels was optional for the participants. The data collection phase lasted one to three days, with valid samples ranging from 52 to 1366 minutes. Sensor data availability varied due to factors such as phone model and usage patterns. After the data collection phase, volunteers were invited to interact with SensorChatC in-person through a chat-based graphical interface. They were encouraged to ask any questions about their lives during the data collection phase and observe SensorChatC’s response generation in real time. Finally, we gathered feedback from the participants including ratings on answer content, latency, and practical value of SensorChat on a scale from 1 to 5. We specifically select SensorChatC for the user study to evaluate SensorChat’s full functionality in real-world conditions. We leave user evaluation and optimization of SensorChatE in a purely edge scenario for future work.
User Feedback Results Fig. 12 displays the quantitative feedback ratings gathered from the eight participants. SensorChat received an average score of 3.12 for answer content, 4.50 for latency, and 4.00 for practical utility, supporting SensorChat’s applicability in real life scenarios. Participant comments praised SensorChat’s natural responses, e.g., “The answers were formatted in an easy to understand way”. However, concerns were also raised about quantitative accuracy, such as “some numbers were a little off” and “it mentioned activities I never did”. We attribute these issues to the limited performance of the sensor encoder when generalizing from a dataset to real-world users, as we observed noisy outputs from the sensor query stage, occasionally detecting activities that never occurred. Fortunately, generalizing models trained on one dataset to different settings is a well-studied problem (Xu et al., 2023b). Integrating these techniques into SensorChat could enhance its practical performance, which we leave for future investigation.
Generalization to Qualitative and Open-Ended Questions Participants have asked creative and open-ended questions to SensorChat that do not exist in SensorQA, providing a more comprehensive assessment of SensorChat’s practical utility. For example, one participant asked, “How would you rate my lifestyle and what improvements do you suggest?” SensorChat responded with, “I would rate this lifestyle as 2 out of 5 stars and would suggest you improve on your exercise and talking,” based on the duration of each activity where exercise and talking were lacking. The three-stage pipeline, especially the explicit sensor data query stage, enables SensorChat to analyze users’ overall lifestyle and fitness, in addition to answering quantitative questions in SensorQA (Reichman et al., 2025). Specifically, SensorChat achieves this through comprehensively querying a list of activity schedule in the sensor data query stage and then analyzing them with LLM during the answer assembly stage. Participants appreciated SensorChat’s ability to effectively handle both quantitative and qualitative questions, e.g., ”I think it could be very useful to be able to answer these qualitative and quantitative questions about my lifestyle,” highlighting its broad applicability to diverse real-life scenarios.
Figure 12. Satisfactory ratings from eight participants on three questions.
8. Discussion & Future Work
Generalizability of SensorChat In the quantitative study, we mainly focus on SensorQA (Reichman et al., 2025) as it is the only available sensor-based QA dataset that aligns with our problem statement. However, SensorChat is not limited to this dataset. As demonstrated in the user study (Sec.7), SensorChat can generalize to broader real-life scenarios and answer high-level qualitative questions. Moreover, the design of SensorChat can be easily expanded. Specifically, SensorChat can incorporate more sensors via adding additional modality encoders. It can also adapt to more diverse questions by expanding the solution templates or using more powerful LLMs such as OpenAI’s o3-mini (ope, 2025) or DeepSeek (dee, 2025) during question decomposition. Last but not least, SensorChat can serve a larger user base by collecting more data and equipping with better generalization designs such as (Xu et al., 2023b). SensorChat is a general and flexible framework designed for fusing long-duration multimodal sensors knowledge with language and answering versatile questions.
Personalization of SensorChat This work emphasizes SensorChat’s ability to generalize across a broad range of questions and users. However, providing personalized responses is equally crucial, particularly for health-related analyses. We foresee the potential to tailor SensorChat for individual users by personalizing the LLMs (Kim et al., 2024), a direction we intend to explore in future research.
Evaluation Metrics SensorChat uses traditional NLP metrics and accuracy to evaluate the quality and precision of generated answers. While these metrics excel in objectiveness, they do not capture subjective user satisfaction, which is crucial for assessing a system’s readiness for a broad market. We leave a comprehensive evaluation of SensorChat with subjective metrics as our future work.
9. Conclusion
Natural interactions between users and multimodal sensors can unlock the full potential of sensor data in real-world applications. However, existing systems struggle to process long-duration, high-dimensional sensor data, often resulting in unsatisfactory answers to long-term questions. In this paper, we present SensorChat, the first edge system designed to handle long-duration, time-series sensor data for both qualitative and quantitative question answering. SensorChat introduces a novel three-stage pipeline including question decomposition, sensor data query and answer assembly. All stages and interfaces are carefully designed to effectively fuse the natural language and the sensor modality. Evaluation results and user studies demonstrate SensorChat’s effectiveness in answering a wide range of qualitative and quantitative questions. The cloud version of SensorChat system supports real-time interactions, while the edge version is optimized to run on Jetson-level devices.
Acknowledgements.
This work was supported in part by National Science Foundation under Grants #2003279, #1826967, #2100237, #2112167, #1911095, #2112665, #2120019, #2211386 and in part by PRISM and CoCoSys, centers in JUMP 2.0, an SRC program sponsored by DARPA.
Appendix A Summarization Function Design
We design the following summarization functions to handle various real-life QA scenarios.
We emphasize that the function set can be expanded in the future to accommodate a broader range of queries. However, since the summarization function is selected by the LLM during question decomposition, it is crucial to limit the number of functions to avoid overwhelming or confusing the model. Fortunately, one strength of SensorChat is its ability to transform questions into existing functions during question decomposition, eliminating the need to create new ones unnecessarily. For instance, a question like ”What did I do right after waking up?” can be decomposed into calls to DetectFirstTime and DetectLastTime, rather than requiring the creation of a new function such as DetectNextActivity. To enable this transformation, we include a list of available functions in the prompt, as shown in Fig. 6(a). This approach ensures that the current set of six summarization functions is sufficient to effectively handle the vast majority of user queries.

### v2

@@ Appendix F Ablation Studies
Table 7. Impact of three major stage in SensorChatC. Bold values highlight the best results. The full SensorChat with all three stages achieve the best results, highlighting the indispensable contribution from each stage.
@@
Impact of Sensor Features and Pretraining Loss Functions We next evaluate the impact of sensor features and loss functions during offline encoder pretraining. Specifically, we compare using statistical features (e.g., mean acceleration) versus raw time series inputs, and IMU2CLIP loss (Moon et al., 2023b) versus our proposed contrastive sensor-text pretraining loss for partial context (see Sec.4.3.1). For IMU2CLIP, text samples are generated by combining all valid labels into one sentence. We report the online querying accuracy and multiple-choice answer accuracy to assess the influence to sensor information extraction. As shown in Table 8, statistical features result in low querying and answer accuracies. This validates our motivation to design SensorChat that high-frequency time series sensor data are critical for fine-grained activity information. IMU2CLIP training, even with fine-grained data, yields poorer querying and answering accuracies, highlighting its limited ability associating sensor embeddings from partial text query. Our proposed loss function, which aligns sensor and text encoders for partial context queries, proves more effective. These findings emphasize the importance of selecting appropriate sensor features and loss functions during pretraining in order to achieve high-performance QA.
@@
Table 9. Impact of various design choices for LLMs in SensorChatC. Bold values highlight the best results. The results highlight the necessity of in-context learning (ICL) and chain-of-thought (CoT) during question decomposition, and using finetuning instead of zero- or few-shot learning during answer assembly.
@@ Appendix G Sensitivity Analyses
Figure 12. Sensitivity of key hyperparameters including learning rate (leftmost), rank in LoRA finetuning (middle left), LLM generating temperature (middle right) and query threshold in SensorChat (right most).
@@
Appendix H Impact of Sensor Data Duration
Figure 13. Performance of SensorChat compared to representative baselines on SensorQA (Reichman et al., 2025), categorized by the duration of sensor data associated with each user question.
Fig.13 presents the performance of SensorChat and representative baselines across different sensor data durations using SensorQA (Reichman et al., 2025). We categorize user questions based on the associated sensor data into single-day and multi-day groups and report short answer accuracy. The baselines include DeepSQA (Xing et al., 2021), LLMSense (Ouyang and Srivastava, 2024), and LLaMA-Adapter (Zhang et al., 2023), each representing a mainstream approach. As discussed in Sec. 6.3, these methods struggle with long-duration, high-frequency sensor data. Fig.13 further shows that their performance declines when scaling from single-day to multi-day inputs: LLaMA-Adapter drops from 0.31 to 0.23, and LLMSense from 0.25 to 0.21. This highlights their limited ability to reason over long-term sensor data. Specifically, LLaMA-Adapter compresses sensor signals into embeddings before passing them to the LLM, which becomes less effective as the time span grows. LLMSense’s drop in performance stems from LLMs’ known difficulty with long-context reasoning (Huang et al., 2023; Gu and Dao, 2023), often resulting in confusion and inaccurate responses. In contrast, SensorChat achieves higher accuracy on multi-day queries (0.58) than on single-day ones (0.46), demonstrating its strength in handling long-duration questions. However, this does not mean SensorChat underperforms on short durations. The slightly lower single-day accuracy is due to the nature of those queries, which demand precise time-based answers more frequently than multi-day queries, such as, “How long did I exercise today?” As discussed in Sec. 6.3, short answer accuracy is a stricter metric for such questions, leading to lower scores compared to multi-day queries (e.g., “On which day did I exercise?”).
Appendix I Generalization and Robustness
@@
Figure 14. Generalization performance of SensorChat ’s answer assembly stage (a) and online query stage (b) when trained on data from different users. Results show that the answer assembly stage generalizes well across users, while the sensor data query stage is more sensitive to shifts in user data distribution.
@@
Generalizability Across Sensor Device Placements We acknowledge that sensor data from vastly different body positions could introduce ambiguity in the encoder. However, detailed device placement information may not always be available, depending on the sensor setup. In this work, SensorChat adopts a generalized approach without assuming prior knowledge of sensor placement. The key idea that SensorChat could work in this case is that, by training on a large-scale dataset collected from diverse users with varying device placements, the model learns a general similarity association between sensor data and context labels. When the user queries about a specific context of their lives, SensorChat could retrieve all sensor recordings sufficiently similar using the using learned similarities. Additionally, if device placement information is available, we foresee that integrating SensorChat with recent sensor context augmentation techniques, such as (Chowdhury et al., 2025), could further enhance performance. We leave this as an avenue for future exploration.
Appendix J Graphical User Interface of SensorChat
Figure 15. A prototype Graphical User Interface (GUI) for SensorChat. The left panel displays an optional activity visualization based on model predictions from raw sensor data. We plan to enhance interactivity by allowing users to select specific days and context labels of interest in the future. The right panel shows the main conversation between users and the SensorChat system: user questions appear in blue box on the right, and SensorChat responses appear in gray box on the left.
Figure 15 presents a prototype of the SensorChat GUI used in our user study. Users primarily interact with SensorChat by posing questions and receiving responses, as illustrated in the right panel of the figure. These questions may be either quantitative or qualitative. In this example, the system operates with a low temperature setting for answer generation (see Appendix G for additional sensitivity experiments), resulting in concise responses that emphasize key facts. This temperature parameter can be adjusted to better align with individual user preferences. The left panel of the GUI includes an optional activity visualizer that provides users with an overview of their schedule. This visualizer can help remind users of their schedule and guide them toward more insightful questions. For lightweight deployments, such as the SensorChatE mobile app on edge devices, this visualizer can be disabled, retaining only the conversation panel for streamlined interaction.
We emphasize that this interface represents an initial prototype of SensorChat. Future enhancements will focus on improving interactivity, such as enabling users to select specific dates and context labels for visualization in the left panel. Additionally, we plan to incorporate a raw sensor data viewer for users interested in inspecting signals captured from their mobile or wearable devices.
Appendix K Questions Asked in The Real User Study
The goal of the real user study is to evaluate SensorChat’s generalizability and practicality in real-world scenarios, extending beyond the predefined scope of SensorQA (Reichman et al., 2025). We notice that some participant questions overlap with those in SensorQA, particularly when they are quantitative. The participants also raised a variety of qualitative questions that reflect their personal interests and are not included in SensorQA. After interacting with SensorChat and posing a mix of both qualitative and quantitative questions, participants reported fairly high satisfaction, demonstrating SensorChat’s adaptability and usefulness in real-life scenarios.
For example, the following quantitative questions are raised by the participants, which are similar to those in SensorQA:
How much time did I spend sitting down vs standing up? How much time did I spend doing lab work? How much time on average did I spend on entertainment for the past 2 days? Did I perform any household chores, if so, what shores did I do? What did I do after having lunch?
The following qualitative questions are posed by the participants during real user study, which are not present in SensorQA:
What do you suggest to make me more active? How would you rate this lifestyle and what do you suggest that I improve on? Did I exercise enough for a young adult? Do I like cycling? Am I attentive when talking to my co-workers?